Find the greatest number of four digits when divided by 3,5,7,9 leaves remainders 1,3,5,7 respectively
1
Expert's answer
2019-05-10T12:53:40-0400

The greatest number of four digits is "9999."


"LCM(3,5,7,9)=5\times7\times9=315"

Dividing "9999" by "315" we find that is largest four digit number that is evenly divisible by 3 , 5 , 7 and 9 is "9765."

Common difference between divisors and respective remainders


"3-1=5-3=7-5=9-7=2"

The required number is


"9765-2=9763"

The greatest number of four digits when divided by 3,5,7,9 leaves remainders 1,3,5,7 respectively is "9763."